# An Extension of the Adiabatic Theorem

### v1

@@ Abstract
We examine the validity of a potential extension of the adiabatic theorem to quantum quenches, i.e., non-adiabatic changes. In particular, the Transverse Field Ising Model (TFIM) and the Axial Next Nearest Neighbour Ising (ANNNI) model are studied. The proposed extension of the adiabatic theorem is stated as follows: Consider the overlap between the initial ground state and the post-quench Hamiltonian eigenstates for quenches within the same phase. This overlap is largest for the post-quench ground state. In the case of the TFIM, this conjecture is confirmed for both the paramagnetic and ferromagnetic phases numerically and analytically. In the ANNNI model, the conjecture could be analytically proven for a special case. Numerical methods were employed to investigate the conjecture’s validity beyond this special case.
@@ I Introduction
Understanding the connection between ground states and eigenstates after a quantum quench can advance our understanding of non-equilibrium dynamics in quantum systems. Specifically, studying how ground states behave after quenches sheds light on energy distributions and phase transition dynamics. Hence, predicting overlaps between eigenstates can provide important insights into the mechanisms governing quench dynamics in non-equilibrium many-body systems.
Quantum quenches and adiabatic time evolution represent two extremes in the study of quantum dynamics, both of which have been the subject of extensive theoretical and numerical research. Previous experimental [1, 2, 3, 4] and theoretical [5, 6, 3] works have explored the evolution of ground states in quenched systems and how memory of the system’s initial state is retained. Quench dynamics have been used to study and identify phase transitions, either directly [7, 8] or through out-of-time-order correlators (OTOCs) [9]. This work investigates the extent to which statements about ground state overlaps hold in the context of quantum quenches. In that sense, it seeks to extend the adiabatic theorem to maximally non-adiabatic changes.
After presenting the conjecture and its connection to the adiabatic theorem in section II we present analytical results for two specific models, the Transverse Field Ising Model (TFIM) and the Axial Next Nearest Neighbour Ising (ANNNI) model. For the TFIM we conduct a full analytical proof in section III, whereas for the ANNNI model only a special case is proven analytically in section IV. To investigate the conjecture’s validity beyond this special case, section V contains a numerical analysis. In section VII we summarise our results and give an outlook to future investigations.
@@ II Conjecture
The adiabatic theorem, formulated by Max Born and Vladimir Fock in 1928, states [10]:
@@
This statement can be re-expressed in terms of quantum state overlaps: consider a quantum system with an initial Hamiltonian $\mathcal{H}_{i}$ at time $t<0$, whose eigenstates are denoted by $|\psi_{n}\rangle$. Without loss of generality, assume the system to be in its ground state, with an excitation gap $\Delta$ above it. At $t=0$, a perturbation acts on the system, such that the eigenstates evolve under the time-dependent Schrödinger equation to $|\psi_{n}(t>0)\rangle$. Let $\mathcal{H}_{f}$ denote the perturbed Hamiltonian, which typically has a different set of eigenstates $|\widetilde{\psi_{n}}\rangle$. For a perturbation with arbitrarily large ramping time $\tau$, the overlap of the time-evolved initial ground state and the final ground state is approaching 1 arbitrarily closely:
@@
Any time-evolution of a quantum system is called “adiabatic” if it happens slowly enough. That is the case if the evolution is much slower than the inverse energy gap size, $\tau\gg\Delta^{-1}$.
@@
[…]
@@
As a necessary condition, we restrict the validity of the conjecture to systems with a continuous spectrum. Since the concept of phase transitions is only well-defined in the thermodynamic limit, the conjecture is expected to hold asymptotically for finite systems.
@@ III Transverse Field Ising Model
$H_{\text{TFIM}}=-J\sum_{j}\left(\sigma_{j}^{z}\sigma_{j+1}^{z}+h\sigma_{j}^{x}\right),$ (3)
where $\sigma^{x}$ and $\sigma^{z}$ denote Pauli matrices. The parameter $J$ denotes an overall energy scale, subsequently set equal to one for convenience, while $h>0$ defines the strength of the transverse magnetic field. At zero temperature, this model undergoes a phase transition from a gapped ferromagnet to a gapped paramagnet at a critical value of $h=J$ [11].
After Jordan-Wigner [12] and Bogoliubov [13] transformations, the Hamiltonian reads
$H_{\text{TFIM}}=\sum_{k>0}\bigg{(}\epsilon_{k}\bigg{(}\gamma_{k}^{\dagger}\gamma_{k}-\gamma_{-k}\gamma_{-k}^{\dagger}\bigg{)}-h\bigg{)}.$ (4)
This is the starting point for the subsequent calculations, where the general idea is to quench from an initial (pre-quench) Hamiltonian $H_{i}$ to a final (post-quench) Hamiltonian $H_{f}$:
$H_{i}\ \xrightarrow{t\ll\Delta^{-1}}\ H_{f}.$ (5)
In general, these Hamiltonians $H_{i}$ and $H_{f}$ have different eigenbases and the pre-quench ground state expressed in the post-quench eigenbasis takes the form
@@
This Eq. (6) is valid for any value of $h$, and therefore applies in both the paramagnetic and the ferromagnetic phase.
@@
$B(k):=-i\frac{V_{k}}{U_{k}}=-i\ \tan(\theta_{k}(h_{i})-\theta_{k}(h_{f})),$ (7)
@@
$\begin{split}U_{k}&=\cos(\theta_{k}(h_{i}))\cos(\theta_{k}(h_{f}))+\sin(\theta_{k}(h_{i}))\sin(\theta_{k}(h_{f})),\\ V_{k}&=\cos(\theta_{k}(h_{i}))\sin(\theta_{k}(h_{f}))-\sin(\theta_{k}(h_{i}))\cos(\theta_{k}(h_{f})).\end{split}$ (8)
For the excited states of the post-quench Hamiltonian in its own eigenbasis, there are three different possibilities:
@@
$|1,0\rangle$ $=\gamma_{-k}^{\dagger}|0,0\rangle,\text{ and}$ (10)
@@
The excitations in Eq.s (9) and (10) involve unpaired creation operators, resulting in zero overlap with the pre-quench ground state. The more interesting excitations are those in Eq. (11), as they involve paired creation operators and can have non-zero overlaps with the pre-quench ground state.
@@
$\begin{split}|\langle\text{GS}(h)\mid 0,0\rangle|^{2}&=\bigg{|}\frac{1}{\mathcal{N}}\langle 0,0\mid\exp{\bigg{(}\sum_{k>0}B(k)\gamma_{-k}\gamma_{k}\bigg{)}}\mid 0,0\rangle\bigg{|}^{2}\\ &=\frac{1}{\mathcal{N}^{2}}\bigg{|}\langle 0,0\mid\prod_{k>0}\bigg{(}1+B(k)\gamma_{-k}\gamma_{k}\bigg{)}\mid 0,0\rangle\bigg{|}^{2}\\ &=\frac{1}{\mathcal{N}^{2}}.\end{split}$ (12)
Then, we compute the overlap of the pre-quench ground state with the post-quench excited state as defined in Eq. (11):
$\begin{split}|\langle\text{GS}(h)\mid 1,1\rangle|^{2}&=\frac{1}{\mathcal{N}^{2}}\bigg{|}\langle 0,0\mid\exp{\bigg{(}\sum_{k>0}B(k)\gamma_{-k}\gamma_{k}\bigg{)}}\gamma_{-k^{\prime}}^{\dagger}\gamma_{k^{\prime}}^{\dagger}\mid 0,0\rangle\bigg{|}^{2}\\ &=\frac{1}{\mathcal{N}^{2}}\bigg{|}B(k^{\prime})\bigg{|}^{2}.\end{split}$ (13)
@@
$\bigg{|}\langle\text{GS}(h)\mid 1,1\rangle\bigg{|}^{2}=\frac{1}{\mathcal{N}^{2}}\tan^{2}\bigg{(}\theta_{k}(h_{i})-\theta_{k}(h_{f})\bigg{)}.$ (14)
@@
Expression (14) must be smaller than the ground state overlap (12) for the conjecture to be correct, hence
$\bigg{|}\langle\text{GS}(h)\mid 1,1\rangle\bigg{|}^{2}\overset{!}{<}\bigg{|}\langle\text{GS}(h)\mid 0,0\rangle\bigg{|}^{2}$ (15)

### v2

# Extension of the adiabatic theorem

@@ Abstract
We examine the validity of a potential extension of the adiabatic theorem to quantum quenches, i.e., nonadiabatic changes. In particular, the transverse field Ising model (TFIM) and the axial next nearest neighbor Ising (ANNNI) model are studied. The proposed extension of the adiabatic theorem is stated as follows: Consider the overlap between the initial ground state and the postquench Hamiltonian eigenstates for quenches within the same phase. This overlap is largest for the postquench ground state. In the case of the TFIM, this conjecture is confirmed for both the paramagnetic and ferromagnetic phases numerically and analytically. In the ANNNI model, the conjecture could be analytically proven for a special case. Numerical methods are employed to investigate the conjecture’s validity beyond this special case.
@@ I Introduction
The quantum adiabatic theorem provides conditions under which a perturbed quantum system remains in its instantaneous eigenstate [1]. Beyond its fundamental conceptual significance, the theorem has been extensively used in condensed matter physics. It provided the basis for the introduction of the Berry phase [2], which plays a central role in the modern description of topological states of matter [3, 4]. Moreover, the adiabatic theorem is fundamental to several static and dynamical methods [5, 6] and has been used to establish numerous formal results [7, 8, 9, 10]. Similarly, understanding the connection between ground states and eigenstates after a quantum quench can advance our understanding of non-equilibrium dynamics in quantum systems. Specifically, studying how ground states behave after quenches sheds light on energy distributions and phase transition dynamics. Hence, predicting overlaps between eigenstates can provide important insights into the mechanisms governing quench dynamics in non-equilibrium many-body systems.
Quantum quenches and adiabatic time evolution represent two extremes in the study of quantum dynamics, both of which have been the subject of extensive theoretical and numerical research. Previous experimental [11, 12, 13, 14] and theoretical [15, 16, 13] works have explored the evolution of ground states in quenched systems and how memory of the system’s initial state is retained. Quench dynamics have been used to study and identify phase transitions, either directly [17, 18] or through out-of-time-order correlators [19]. This work investigates the extent to which statements about ground state overlaps hold in the context of quantum quenches. In that sense, it seeks to extend the adiabatic theorem to maximally nonadiabatic changes.
After presenting the conjecture and its connection to the adiabatic theorem in Sec. II we present analytical results for two specific models, the transverse field Ising model (TFIM) and the axial next nearest neighbor Ising (ANNNI) model. For the TFIM we conduct a full analytical proof in Sec. III, whereas for the ANNNI model only a special case is proven analytically in Sec. IV. To investigate the conjecture’s validity beyond this special case, section V contains a numerical analysis. In Sec. VII we summarize our results and provide an outlook on future investigations.
@@ II Conjecture
The adiabatic theorem, formulated by Born and Fock in 1928, states [1],
@@
This statement can be reexpressed in terms of quantum state overlaps: Consider a quantum system with an initial Hamiltonian $\mathcal{H}_{i}$ at time $t<0$, whose eigenstates are denoted by $|\psi_{n}\rangle$. Without loss of generality, assume the system is in its ground state, with an excitation gap $\delta E$ above it. At $t=0$, a perturbation acts on the system, such that the eigenstates evolve under the time-dependent Schrödinger equation to $|\psi_{n}(t>0)\rangle$. Let $\mathcal{H}_{f}$ denote the perturbed Hamiltonian, which typically has a different set of eigenstates $|\widetilde{\psi_{n}}\rangle$. For a perturbation with arbitrarily large ramping time $\tau$, the overlap of the time-evolved initial ground state and the final ground state approaches 1 arbitrarily closely:
@@
Any time evolution of a quantum system is called “adiabatic” if it happens slowly enough. That is the case if the evolution is much slower than the inverse energy gap size, $\tau\gg(\delta E)^{-1}$.
@@
The overlap of the initial ground state $|\text{GS}\rangle$ with any postquench eigenstate $|\widetilde{\psi_{n}}\rangle$ is maximal if $|\widetilde{\psi_{n}}\rangle$ is the postquench ground state, as long as both Hamiltonians are in the same phase:
@@
As a necessary condition, we restrict the validity of the conjecture to systems with a continuous spectrum. The notion of equal phases that we impose for this conjecture is the existence of a continuous path in the space of Hamiltonians with the same symmetries connecting the pre- and postquench Hamiltonians without any gap-closing. In other words, the system remains gapped at every point along this path. Since the concept of phase transitions is well-defined only in the continuous spectrum limit, the conjecture is expected to hold asymptotically for finite systems. The overlap in the conjecture above is closely related to the information-theoretic ground state fidelity [20, 21]
$F(a,\delta a)$ $=|\left\langle\text{GS}(a)\mid\text{GS}(a+\delta a)\right\rangle|$ (3)
$=1-\frac{1}{2}\chi_{F}(a)(\delta a)^{2}+\mathcal{O}((\delta a)^{3}),$
quantifying the resemblance of two ground states. Here, $a$ denotes a parameter of the Hamiltonian and $\delta a$ denotes a small change in $a$. For the thermodynamic limit one defines its intrinsic version as $\mathcal{F}(a,\delta a)=\lim_{N\rightarrow\infty}\frac{1}{N}\ln F(a,\delta a)$ [22, 23, 24] where NN denotes the system size. An important quantity arising in the discussion of ground state fidelities is the fidelity susceptibility $\chi_{F}(a)$, which is a measure of how rapidly the ground state changes under perturbations of a control parameter. If the fidelity susceptibility per lattice site is an analytic function, a finite parameter region in which the ground state overlap remains the largest must exist. Away from quantum critical points and (quantum) phase transitions the fidelity susceptibility is expected to remain an analytic function of the control parameter. It is widely accepted in the community––and supported by much numerical evidence––that the fidelity susceptibility can diverge only in the vicinity of phase transitions [25, 26] and quantum critical points [27, 28]. However, there is no mathematically rigorous proof for it. Because even this weaker statement about the fidelity susceptibility has not been proven (yet), one should not expect to establish a proof for the more general conjecture formulated here. Currently, the only possible approach to verify or disprove the conjecture is with numerical studies, just as is the case for the fidelity susceptibility. Nevertheless, we believe that our conjecture holds in a finite parameter region sufficiently far from phase transitions and quantum critical points.
@@ III Transverse Field Ising Model
$H_{\mathrm{TFIM}}=-J\sum_{j}\left(\sigma_{j}^{z}\sigma_{j+1}^{z}+h\sigma_{j}^{x}\right),$ (4)
where $\sigma^{x}$ and $\sigma^{z}$ denote Pauli matrices. The parameter $J$ denotes an overall energy scale, subsequently set equal to 1 for convenience, while $h>0$ defines the strength of the transverse magnetic field. At zero temperature, this model undergoes a phase transition from a gapped ferromagnet to a gapped paramagnet at a critical value of $h=J$ [29].
After Jordan–Wigner [30] and Bogoliubov [31] transformations, the Hamiltonian in momentum space reads
$H_{\mathrm{TFIM}}=\sum_{k>0}\big[\epsilon_{k}\big(\gamma_{k}^{\dagger}\gamma_{k}-\gamma_{-k}\gamma_{-k}^{\dagger}\big)-h\big],$ (5)
where $\epsilon_{k}=\sqrt{h^{2}-2h\cos(k)+1}$ and $\gamma_{k}^{\dagger}$ and $\gamma_{k}$ denote the Bogoliubov operators
$\gamma_{k}$ $=\cos(\theta_{k})c_{k}+i\sin(\theta_{k})c_{-k}^{\dagger},$ (6)
$\gamma_{k}^{\dagger}$ $=\cos(\theta_{k})c_{k}^{\dagger}-i\sin(\theta_{k})c_{-k},$ (7)
which are in turn defined via the Fourier-transformed operators of a rotated Jordan–Wigner transformation
$\sigma_{i}^{x}$ $=1-2c_{i}^{\dagger}c_{i},$ (8)
$\sigma_{i}^{z}$ $=-\prod_{j<i}\big(1-2c_{j}^{\dagger}c_{j}\big)\big(c_{i}+c_{i}^{\dagger}\big),$ (9)
using the Fourier convention
$c_{k}=\frac{1}{\sqrt{L}}\sum_{j}e^{-ikj}c_{j},\quad c_{k}^{\dagger}=\frac{1}{\sqrt{L}}\sum_{j}e^{ikj}c_{j}^{\dagger}.$ (10)
This is the starting point for the subsequent calculations, where the general idea is to quench from an initial (prequench) Hamiltonian $H_{i}$ to a final (postquench) Hamiltonian $H_{f}$:
$H_{i}\ \xrightarrow{t\ll(\delta E)^{-1}}\ H_{f}.$ (11)
In general, these Hamiltonians $H_{i}$ and $H_{f}$ have different eigenbases and the prequench ground state expressed in the postquench eigenbasis takes the form
@@
$B(k):=-i\frac{V_{k}}{U_{k}}=-i\tan[\theta_{k}(h_{i})-\theta_{k}(h_{f})],$ (13)
@@
$U_{k}=\cos[\theta_{k}(h_{i})]\cos[\theta_{k}(h_{f})]+\sin[\theta_{k}(h_{i})]\sin[\theta_{k}(h_{f})],$
$V_{k}=\cos[\theta_{k}(h_{i})]\sin[\theta_{k}(h_{f})]-\sin[\theta_{k}(h_{i})]\cos[\theta_{k}(h_{f})].$ (14)
These expressions and the ground state (12) are valid for any value of $h$, and therefore apply in both the paramagnetic and ferromagnetic phase. The angles $\theta_{k}(h)$ are given by the parametrization of the Bogoliubov transformation [29]
$\tan(2\theta_{k})=\frac{\sin k}{h-\cos k}\ \Leftrightarrow\ \theta_{k}(h)=\frac{1}{2}\arctan\left(\frac{\sin k}{h-\cos k}\right)$ (15)
and are limited to the interval $0,\frac{\pi}{2}]$. As $k=2\pi\frac{n}{L}$, with $n\in\mathbb{N}$ in the interval $1,\frac{L}{2}]$ and $L$ being the spin chain length, the values for $k$ cover half a period from 0 to $\pi$.
For the excited states of the postquench Hamiltonian in its own eigenbasis, there are three different possibilities:
@@
$|1,0\rangle$ $=\gamma_{-k}^{\dagger}|0,0\rangle,$ (17)
@@
The excitations in Eqs. (16) and (17) involve unpaired creation operators, resulting in zero overlap with the prequench ground state. The more interesting excitations are those in Eq. (18), as they involve paired creation operators and can have non-zero overlaps with the prequench ground state.
@@
$|\langle\text{GS}(h)\mid 0,0\rangle|^{2}$ $=\bigg|\frac{1}{\mathcal{N}}\langle 0,0\mid\exp{\left(\sum_{k>0}B(k)\gamma_{-k}\gamma_{k}\right)}\mid 0,0\rangle\bigg|^{2}$
$=\frac{1}{\mathcal{N}^{2}}\bigg|\langle 0,0\mid\prod_{k>0}\bigg(1+B(k)\gamma_{-k}\gamma_{k}\bigg)\mid 0,0\rangle\bigg|^{2}$
$=\frac{1}{\mathcal{N}^{2}}.$ (19)
Then, we compute the overlap of the prequench ground state with the postquench excited state as defined in Eq. (18):
$\Big|\langle\text{GS}$ $(h)\mid 1,1\rangle\Big|^{2}$
$=\frac{1}{\mathcal{N}^{2}}\left|\langle 0,0\mid\exp{\left(\sum_{k>0}B(k)\gamma_{-k}\gamma_{k}\right)}\gamma_{-k^{\prime}}^{\dagger}\gamma_{k^{\prime}}^{\dagger}\mid 0,0\rangle\right|^{2}$
$=\frac{1}{\mathcal{N}^{2}}\Big|B(k^{\prime})\Big|^{2}.$ (20)
@@
$\bigg|\langle\text{GS}(h)\mid 1,1\rangle\bigg|^{2}=\frac{1}{\mathcal{N}^{2}}\tan^{2}\left[\theta_{k}(h_{i})-\theta_{k}(h_{f})\right].$ (21)
@@
Expression (21) must be smaller than the ground state overlap (III) for the conjecture to be correct; hence
$\Big|\langle\text{GS}(h)\mid 1,1\rangle\Big|^{2}\overset{!}{<}\Big|\langle\text{GS}(h)\mid 0,0\rangle\Big|^{2}$ (22)
$\Leftrightarrow\frac{1}{\mathcal{N}^{2}}\tan^{2}\left[\theta_{k}(h_{i})-\theta_{k}(h_{f})\right]\overset{!}{<}\frac{1}{\mathcal{N}^{2}}$ (23)
$\Leftrightarrow\tan^{2}\left[\theta_{k}(h_{i})-\theta_{k}(h_{f})\right]\overset{!}{<}1.$ (24)
Condition (24) is fulfilled if and only if the argument of $\tan^{2}(x)$ lies in the interval $\frac{\pi}{4},\frac{\pi}{4}]$. We conclude that
$\big|\theta_{k}(h_{i})-\theta_{k}(h_{f})\big|\overset{!}{<}\frac{\pi}{4}.$ (25)
Taking the parametrization of $\theta_{k}(h)$ that was given in Eq. (15) and using the inverse tangent function
$\tan^{-1}(2\theta_{k})=\begin{cases}\arctan(2\theta_{k}),\qquad\ \,\ 2\theta_{k}>0,\\ \arctan(2\theta_{k})+\pi,\quad 2\theta_{k}<0,\end{cases}$ (26)
imply, for the definition of $\theta_{k}(h)$,
[…]
  1. (1)
    Both Hamiltonians lie in the paramagnetic phase. Without loss of generality, it is assumed that $1<h_{i}<h_{f}$}. In the paramagnetic phase, $\theta_{k}=\frac{1}{2}\arctan(\frac{\sin k}{h-\cos k})\ \forall\ h_{i},h_{f},k$} because $\frac{\sin k}{h-\cos k}>0$} for all values of $h_{i},h_{f}\in(1,\infty)$} and $k\in[0,\pi]$}.
    Here $\arctan(\frac{\sin k}{h_{i}-\cos k})$ is a monotonically decreasing function in $h$ for all $k\in[0,\pi]$. Therefore, the maximal difference of $\theta_{k}(h_{i})$ and $\theta_{k}(h_{f})$ is reached for maximally different $h_{i}$ and $h_{f}$. We choose $h_{i}$ to be $1+\epsilon$ with $\epsilon\rightarrow 0$ and $h_{f}$ to be $x\rightarrow\infty$ and obtain
    $\max_{h_{i},h_{f},k}\left(\left|\theta_{k}(h_{i})-\theta_{k}(h_{f})\right|\right)=\frac{\pi}{4}.$ (28)
    As the function approaches the limit from below, conjecture (2) is fulfilled for all possible parameter combinations of $h_{i}$, $h_{f}$ and $k$ in the paramagnetic phase.
  2. (2)
    Both Hamiltonians lie in the ferromagnetic phase. Without loss of generality, it is assumed that $0\leq h_{i}<h_{f}<1$}. For $h\in[0,1)$}, we rewrite Eq. (27) as
    $\theta_{k}(h)=\begin{cases}\frac{1}{2}\arctan(\frac{\sin k}{h-\cos k}),\qquad\quad\ \ k>\arccos(h),\\ \frac{1}{2}\big[\arctan(\frac{\sin k}{h-\cos k})+\pi\big],\quad k<\arccos(h).\end{cases}$ (29)
    We consider $\arccos(h)<k=\arccos(h)+\nu$} with $\nu>0$:
    $\quad\max_{h_{i},h_{f},k}\left[\left|\theta_{k}(h_{i})-\theta_{k}(h_{f})\right|\right]$ (30)
    $=\max_{h_{i},h_{f},\nu}\bigg[\frac{1}{2}\bigg|\arctan\left(\frac{\sin[\arccos(h_{i})+\nu]}{h_{i}-\cos[\arccos(h_{i})+\nu]}\right)$
    $\quad\quad\quad\ -\arctan\left(\frac{\sin[\arccos(h_{f})+\nu]}{h_{f}-\cos[\arccos(h_{f})+\nu]}\right)\bigg|\bigg].$
    As $\arctan\left(\cdot\right)$ is a monotone function in $h$ for all $\nu\in(0,\pi]$, the maximal difference of $\theta_{k}(h_{i})$ and $\theta_{k}(h_{f})$ is reached for maximally different $h_{i}$ and $h_{f}$. Thus, we choose $h_{i}$ to be zero and $h_{f}$ to be $1-\epsilon$ with $\epsilon\rightarrow 0$, such that
    $\max_{h_{i},h_{f},k}\left[\left|\theta_{k}(h_{i})-\theta_{k}(h_{f})\right|\right]$
    $=\max_{\nu}\bigg[\frac{1}{2}\bigg|\arctan\left(\frac{\sin(\frac{\pi}{2}+\nu)}{\cos(\frac{\pi}{2}+\nu)}\right)$
    $\quad\quad\quad+\arctan\left(\frac{\sin(\nu)}{1-\cos(\nu)}\right)\bigg|\bigg]$ (31)
    $\max_{h_{i},h_{f},k}\left[\left|\theta_{k}(h_{i})-\theta_{k}(h_{f})\right|\right]=\frac{\pi}{8}<\frac{\pi}{4}.$ (32)
  3. (3)
    Again, both Hamiltonians lie in the ferromagnetic phase. As before, without loss of generality, it is assumed that $0<h_{i}<h_{f}<1$. However, now we consider the case of $\arccos(h)>k=\arccos(h)-\nu$ with $\nu>0$. From $k\in[0,\pi]$ it follows that $\nu\overset{!}{\in}[-\frac{\pi}{2},0)$.
    $\max_{h_{i},h_{f},k}\left[\left|\theta_{k}(h_{i})-\theta_{k}(h_{f})\right|\right]$ (33)
    $=\max_{h_{i},h_{f},\nu}\bigg[\frac{1}{2}\bigg|\arctan\left(\frac{\sin[\arccos(h_{i})-\nu]}{h_{i}-\cos[\arccos(h_{i})-\nu]}\right)+\pi$
    $\qquad\qquad\qquad\qquad-\arctan\left(\frac{\sin[\arccos(h_{f})-\nu]}{h_{f}-\cos[\arccos(h_{f})-\nu]}\right)+\pi\bigg|\bigg].$
    As $\arctan(\cdot)$ is a monotone function in $h$ for all $\nu\in[-\frac{\pi}{2},0)$, the maximal difference of $\theta_{k}(h_{i})$ and $\theta_{k}(h_{f})$ is again reached for maximally different $h_{i}$ and $h_{f}$. Thus we again choose $h_{i}$ to be zero and $h_{f}$ to be $1-\epsilon$ with $\epsilon\rightarrow 0$, such that
    $\max_{h_{i},h_{f},k}\left[\left|\theta_{k}(h_{i})-\theta_{k}(h_{f})\right|\right]=\frac{\pi}{8}<\frac{\pi}{4},$ (34)
    and conjecture (2) is confirmed for this second choice of $\theta_{k}(h)$.
One can easily check that condition (24) is not necessarily satisfied once the phase transition is crossed. In fact, in the vicinity of the critical point at $h=1$, the condition is highly sensitive to even small changes in $h$. For instance, choosing $h_{i}<h_{f}$} (without loss of generality) in different phases, with $h_{i}=1-\epsilon$} and $h_{f}=1+\epsilon$}, an $\epsilon$ of order $10^{-2}$ is already sufficient to violate condition (24) for some $k$. As $\epsilon$ increases further, the set of momenta for which the condition fails grows correspondingly.
However, they have an even smaller overlap with the prequench ground state than the excitation (18), as each excitation pair contributes another factor further diminishing the overlap with the prequench ground state.
IV Axial Next Nearest Neighbor Ising Model
Figure 1: Phase diagram of the ANNNI model. The red line within the paramagnetic phase indicates the Peschel–Emery line. Adapted from [32].
The second model investigated in this work is a modification of the transverse field Ising model, incorporating additional next nearest neighbor interactions. It is known as the anisotropic or axial next nearest neighbor Ising model and has been studied extensively in the last few decades [33, 34, 32, 17, 18]. Alongside the nearest neighbor coupling $J$ and the coupling $h$ to an external magnetic field, the interaction between next nearest neighbors can be tuned by a parameter $\Delta$ in the Hamiltonian [32]